### v1

# Entanglement witnesses and separability criteria based on generalized equiangular tight frames

Abstract
We use operators from generalized equiangular measurements to construct positive maps. Their positivity follows from the inequality for indices of coincidence corresponding to few equiangular tight frames. These maps give rise to entanglement witnesses, which include as special cases many important classes considered in the literature. Additionally, we introduce separability criteria based on the correlation matrix and analyze them for various types of measurements.
1 Measurements from generalized equiangular tight frames
A set $\{P_{k}:\,k=1,\ldots,M\}$ of rank-1 projectors $P_{k}$ acting on the Hilbert space $\mathcal{H}\simeq\mathbb{C}^{d}$ is an equiangular tight frame if $\mathrm{Tr}(P_{k}P_{\ell})=c\mathrm{Tr}(P_{k})\mathrm{Tr}(P_{\ell})$ for all $k\neq\ell$ and $\sum_{k=1}^{M}P_{k}=\gamma\mathbb{I}_{d}$ with $\gamma>0$ Strohmer . If $P_{k}$ are no longer rank-1, then they form a generalized equiangular tight frame GEAM . Now, take collections of $N$ generalized equiangular tight frames $\mathcal{P}_{\alpha}=\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha}\}$, $\alpha=1,\ldots,N$, whose elements sum up to $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$, where $\gamma_{\alpha}$ is a probability distribution. Additionally, assume that they are complementary PetzRuppert , which means that $\mathrm{Tr}(P_{\alpha,k}P_{\beta,\ell})=f\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\beta,\ell})$, $\alpha\neq\beta$. Under these conditions, one defines the corresponding positive, operator-valued measure (POVM).
Definition 1.
(GEAM ) A collection of $N$ generalized equiangular lines $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ satisfying $\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}=\gamma_{\alpha}\mathbb{I}_{d}$ for a probability distribution $\gamma_{\alpha}$ is a generalized equiangular measurement provided that
$\begin{split}\mathrm{Tr}(P_{\alpha,k})&=a_{\alpha},\\ \mathrm{Tr}(P_{\alpha,k}^{2})&=b_{\alpha}\mathrm{Tr}(P_{\alpha,k})^{2},\\ \mathrm{Tr}(P_{\alpha,k}P_{\alpha,\ell})&=c_{\alpha}\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\alpha,\ell}),\qquad k\neq\ell,\\ \mathrm{Tr}(P_{\alpha,k}P_{\beta,\ell})&=f\mathrm{Tr}(P_{\alpha,k})\mathrm{Tr}(P_{\beta,\ell}),\qquad\alpha\neq\beta,\end{split}$ (1)
where
$a_{\alpha}=\frac{d\gamma_{\alpha}}{M_{\alpha}},\qquad c_{\alpha}=\frac{M_{\alpha}-db_{\alpha}}{d(M_{\alpha}-1)},\qquad f=\frac{1}{d},$ (2)
$\frac{1}{d}<b_{\alpha}\leq\frac{1}{d}\min\{d,M_{\alpha}\}.$ (3)
Generalized equiangular measurements can be constructed from Hermitian orthonormal bases $\{G_{0}=\mathbb{I}_{d}/\sqrt{d},G_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ with traceless $G_{\alpha,k}$. Namely, one has GEAM
$P_{\alpha,k}=\frac{a_{\alpha}}{d}\mathbb{I}_{d}+\tau_{\alpha}H_{\alpha,k},$ (4)
where the parameter
$\tau_{\alpha}^{2}=\frac{a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})}{M_{\alpha}(\sqrt{M_{\alpha}}+1)^{2}}$ (5)
is chosen this way to guarantee $P_{\alpha,k}\geq 0$. The traceless operators $H_{\alpha,k}$ are related to $G_{\alpha,k}$ in the following way,
$H_{\alpha,k}=\left\{\begin{aligned} &G_{\alpha}-\sqrt{M_{\alpha}}(1+\sqrt{M_{\alpha}})G_{\alpha,k},\quad k=1,\ldots,M_{\alpha}-1,\\ &(1+\sqrt{M_{\alpha}})G_{\alpha},\qquad k=M_{\alpha},\end{aligned}\right.$ (6)
where $G_{\alpha}=\sum_{k=1}^{M_{\alpha}-1}G_{\alpha,k}$.
There exists an important class of generalized equiangular measurements that are conical 2-designs Graydon ; Graydon2 – that is, they satisfy
$\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}P_{\alpha,k}\otimes P_{\alpha,k}=\kappa_{+}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\kappa_{-}\mathbb{F}_{d}$ (7)
for real parameters $\kappa_{+}\geq\kappa_{-}>0$, where $\mathbb{F}_{d}=\sum_{m,n=1}^{d}|m\rangle\langle n|\otimes|n\rangle\langle m|$ is the flip operator. It has been shown in ref. GEAM that a linear dependence between purity $\mathrm{Tr}(\rho^{2})$ of a state $\rho$ and the corresponding index of coincidence Rastegin5
$\mathcal{C}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2},\qquad p_{\alpha,k}=\mathrm{Tr}(P_{\alpha,k}\rho),$ (8)
exists if and only if $P_{\alpha,k}$ form a conical 2-design.
Proposition 1.
The generalized equiangular measurement whose parameters satisfy $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, where
$0<S\leq\min\left\{\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}},\frac{d-1}{M_{\alpha}-1}\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}\right\},$ (9)
is a conical 2-design with real parameters
$\kappa_{+}=\mu-\frac{S}{d},\qquad\kappa_{-}=S,\qquad\mu=\frac{1}{d}\sum_{\alpha=1}^{N}a_{\alpha}\gamma_{\alpha}.$ (10)
Observe that each generalized equiangular tight frame has its own index of coincidence (after rescaling)
$\mathcal{C}_{\alpha}=\frac{1}{\gamma_{\alpha}^{2}}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2},\qquad\frac{1}{\gamma_{\alpha}}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}=1.$ (11)
Therefore, it is possible to derive an inequality for sums of ssquared probabilities, similar to that for few mutually unbiased bases Wu .
Proposition 2.
If $\{P_{\alpha,k}:\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ is a conical 2-design, then
$\mathcal{C}_{L}=\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq S\left(\mathrm{Tr}\rho^{2}-\frac{1}{d}\right)+\mu_{L},$ (12)
where $\mu_{L}=(1/d)\sum_{\alpha=1}^{L}a_{\alpha}\gamma_{\alpha}$ and $1\leq L\leq N$. The equality is reached for $L=N$.
Proof.
In this proof, we follow the method introduced in ref. Rastegin2 . Any quantum state $\rho$ can be represented as
$\rho=\frac{1}{d}\mathbb{I}_{d}+\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}H_{\alpha,k}$ (13)
with real-valued parameters $r_{\alpha,k}$. Using the properties of the Hermitian operators $H_{\alpha,k}$,
$\begin{split}\mathrm{Tr}(H_{\alpha,k}^{2})&=(M_{\alpha}-1)(\sqrt{M_{\alpha}}+1)^{2},\\ \mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})&=-(\sqrt{M_{\alpha}}+1)^{2},\qquad\ell\neq k,\\ \mathrm{Tr}(H_{\alpha,k}H_{\beta,\ell})&=0,\qquad\beta\neq\alpha.\end{split}$ (14)
we calculate
$p_{\alpha,k}=\mathrm{Tr}(\rho P_{\alpha,k})=\frac{a_{\alpha}}{d}+\tau_{\alpha}\mathrm{Tr}(\rho H_{\alpha,k})=\frac{a_{\alpha}}{d}+\tau_{\alpha}\sum_{\ell=1}^{M_{\alpha}}r_{\alpha,\ell}\mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})=\frac{a_{\alpha}}{d}+\tau(\sqrt{M_{\alpha}}+1)^{2}(M_{\alpha}r_{\alpha,k}-r_{\alpha}),$ (15)
where $r_{\alpha}=\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}$. Now, the sum of squares
$\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}}+\tau_{\alpha}^{2}M_{\alpha}(\sqrt{M_{\alpha}}+1)^{4}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).$ (16)
Remembering that $\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}=\gamma_{\alpha}$, one finds
$\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}=\frac{a_{\alpha}^{2}M_{\alpha}}{d^{2}}\leq\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq\gamma_{\alpha}^{2},$ (17)
which shows that the second term in eq. (16) is positive. Now, we compute the sum over an incomplete set of indices $\alpha=1,\ldots,L$;
$\begin{split}\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}&=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+\sum_{\alpha=1}^{L}\tau_{\alpha}^{2}M_{\alpha}(\sqrt{M_{\alpha}}+1)^{4}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+\sum_{\alpha=1}^{L}a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).\end{split}$ (18)
On the other hand, note that
$\sum_{k,\ell=1}^{M_{\alpha}}r_{\alpha,k}r_{\alpha,\ell}\mathrm{Tr}(H_{\alpha,k}H_{\alpha,\ell})=(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right),$ (19)
and hence the purity of the state $\rho$ is equal to
$\mathrm{Tr}(\rho^{2})=\frac{1}{d}+\sum_{\alpha=1}^{N}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right).$ (20)
After comparing eqs. (18) and (20), it becomes evident that the relation between the sum of squared $p_{\alpha,k}$ and the purity of $\rho$ is linear provided that $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$, which is the condition for a conical 2-design. In this case,
$\begin{split}\sum_{\alpha=1}^{L}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}&=\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+S\sum_{\alpha=1}^{L}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &\leq\frac{1}{d^{2}}\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}+S\sum_{\alpha=1}^{N}(\sqrt{M_{\alpha}}+1)^{2}\left(M_{\alpha}\sum_{k=1}^{M_{\alpha}}r_{\alpha,k}^{2}-r_{\alpha}^{2}\right)\\ &=S\left[\mathrm{Tr}(\rho^{2})-\frac{1}{d}\right]+\mu_{L},\end{split}$ (21)
where $\mu_{L}=\sum_{\alpha=1}^{L}a_{\alpha}^{2}M_{\alpha}/d^{2}=(1/d)\sum_{\alpha=1}^{L}a_{\alpha}\gamma_{\alpha}$. For $L=N$, this result reproduces the index of coincidence $\mathcal{C}$, for which the equality holds GEAM . ∎
Observe that eq. (12) is upper bounded by its value on pure states,
$\mathcal{C}_{L}\leq\widetilde{\mathcal{C}}_{L}=\frac{d-1}{d}S+\mu_{L}.$ (22)
2 Positive maps and entanglement witnesses
Using generalized equiangular measurements that are conical 2-designs, we construct $N$ maps
$\Phi_{\alpha}[X]=\sum_{k,\ell=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}P_{\alpha,k}\mathrm{Tr}(XP_{\alpha,\ell}),$ (23)
where $\mathcal{O}^{(\alpha)}$ are orthogonal rotation matrices preserving the maximally mixed vector $\mathbf{n}_{\ast}=(1,\ldots,1)/\sqrt{d}$. Note that
$\mathrm{Tr}(\Phi_{\alpha}[X])=a_{\alpha}\gamma_{\alpha}\mathrm{Tr}(X),$ (24)
which means that $\Phi_{\alpha}$ are not trace preserving. This is important in proving the following proposition, where trace-preserving $\Phi_{\alpha}$ would impose additional conditions on $P_{\alpha,k}$.
Proposition 3.
The linear map
$\Phi=A\Phi_{0}+\sum_{\alpha=L+1}^{K}\Phi_{\alpha}-\sum_{\alpha=1}^{K}\Phi_{\alpha}$ (25)
with $A=(d-1)S-d(\mu_{K}-2\mu_{L})$ and $\Phi_{0}[X]=\mathbb{I}_{d}\mathrm{Tr}(X)/d$ is positive, where $1\leq L\leq K\leq N$.
Proof.
The condition for positivity of $\Phi$ follows directly from Mehta’s theorem Mehta
$\frac{\mathrm{Tr}(\Phi[P])^{2}}{[\mathrm{Tr}(\Phi[P])]^{2}}\leq\frac{1}{d-1}.$ (26)
It is enough to prove the above inequality for rank-1 projectors $P$ MUBs . The denominator simply reads
$[\mathrm{Tr}(\Phi[P])]^{2}=(A+d\mu_{K}-2d\mu_{L})^{2}.$ (27)
Computing the numerator requires more attention. First, note that
$\begin{split}\mathrm{Tr}(\Phi[P])^{2}=\mathrm{Tr}\Bigg{\{}&A^{2}\Phi_{0}[P]^{2}+\sum_{\alpha,\beta=1}^{L}\Phi_{\alpha}[P]\Phi_{\beta}[P]+\sum_{\alpha,\beta=L+1}^{K}\Phi_{\alpha}[P]\Phi_{\beta}[P]+2A\sum_{\alpha=L+1}^{K}\Phi_{0}[P]\Phi_{\alpha}[P]\\ &-2A\sum_{\alpha=1}^{L}\Phi_{0}[P]\Phi_{\alpha}[P]-2\sum_{\alpha=L+1}^{K}\sum_{\beta=1}^{L}\Phi_{\alpha}[P]\Phi_{\beta}[P]\Bigg{\}}.\end{split}$ (28)
Next, using the properties of the rotation matrix
$\sum_{k=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}=\sum_{\ell=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}=1,\qquad\sum_{k=1}^{M_{\alpha}}\mathcal{O}^{(\alpha)}_{k\ell}\mathcal{O}^{(\alpha)}_{km}=\delta_{\ell m},$ (29)
as well as the properties of $\Phi_{\alpha}$,
$\mathrm{Tr}(\Phi_{0}[P]^{2})=\frac{1}{d},\qquad\mathrm{Tr}(\Phi_{0}[P]\Phi_{\alpha}[P])=\frac{a_{\alpha}\gamma_{\alpha}}{d},\qquad\mathrm{Tr}(\Phi_{\alpha}[P]\Phi_{\beta}[P])=\frac{a_{\alpha}a_{\beta}\gamma_{\alpha}\gamma_{\beta}}{d},\quad\alpha\neq\beta,$ (30)
we simplify eq. (28) to
$\begin{split}d\mathrm{Tr}(\Phi[P])^{2}=(A+d\mu_{K}-2d\mu_{L})^{2}-\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}+d\sum_{\alpha=1}^{K}\mathrm{Tr}(\Phi_{\alpha}[P])^{2},\end{split}$ (31)
where $\mu_{L}=\sum_{\alpha=1}^{L}a_{\alpha}\gamma_{\alpha}/d$, just as defined in Proposition 1. Now, observe that
$\mathrm{Tr}(\Phi_{\alpha}[P]^{2})=a_{\alpha}^{2}\gamma_{\alpha}^{2}c_{\alpha}+a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}$ (32)
with $p_{\alpha,k}=\mathrm{Tr}(PP_{\alpha,k})$ being calculated on a pure state $P$. In this step, we assume that $P_{\alpha,k}$ form a conical 2-design (i.e., $a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=S$) to apply the results from eq. (22). For pure states $\rho=P$, the purity $\mathrm{Tr}(\rho^{2})=1$, and hence
$\sum_{\alpha=1}^{K}\mathrm{Tr}(\Phi_{\alpha}[P]^{2})=\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}c_{\alpha}+S\sum_{\alpha=1}^{K}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{2}\leq\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}c_{\alpha}+\frac{d-1}{d}S^{2}+S\mu_{K}.$ (33)
This way, eq. (34) reduces to
$d\mathrm{Tr}(\Phi[P])^{2}\leq(A+d\mu_{K}-2d\mu_{L})^{2}-\sum_{\alpha=1}^{K}a_{\alpha}^{2}\gamma_{\alpha}^{2}(1-dc_{\alpha})+S[(d-1)S+d\mu_{K}].$ (34)
Finally, note that
$S=a_{\alpha}^{2}(b_{\alpha}-c_{\alpha})=a_{\alpha}\gamma_{\alpha}(1-dc_{\alpha}),$ (35)
and therefore
$d\mathrm{Tr}(\Phi[P])^{2}\leq(A+d\mu_{K}-2d\mu_{L})^{2}+(d-1)S^{2}.$ (36)
Remembering that $A=(d-1)S-d(\mu_{K}-2\mu_{L})$, we arrive at
$d\mathrm{Tr}(\Phi[P])^{2}\leq d(d-1)S^{2}=\frac{d[\mathrm{Tr}(\Phi[P])]^{2}}{d-1},$ (37)
which proves that $\Phi$ is indeed positive. ∎
Corollary 1.
Actually, to define the positive map $\Phi$, it is enough to know $\widetilde{C}_{K}$ and $\widetilde{C}_{L}$ due to
$A=d(2\widetilde{C}_{L}-\widetilde{C}_{K}).$ (38)
The Choi-Jamiołkowski isomorphism Choi ; Jamiolkowski
$W=\sum_{k,\ell=0}^{d-1}|k\rangle\langle\ell|\otimes\Phi[|k\rangle\langle\ell|]$ (39)
establishes a one-to-one correspondence between a positive map $\Phi$ and an entanglement witness $W$, using an operational basis $|k\rangle$ in $\mathbb{C}^{d}$. The witness corresponding to $\Phi$ from Proposition 2 reads
$W=\frac{A}{d}\mathbb{I}_{d}\otimes\mathbb{I}_{d}+\sum_{\alpha=L+1}^{K}J_{\alpha}-\sum_{\alpha=1}^{K}J_{\alpha}$ (40)
where
$J_{\alpha}=\sum_{k,\ell=1}^{M_{\alpha}}\mathcal{O}_{k\ell}^{(\alpha)}\overline{P}_{\alpha,\ell}\otimes P_{\alpha,k}.$ (41)
3 Separability criteria
Take a bipartite composite system with subsystems labeled by $A$ and $B$. On each subsystem, define a conical 2-design that is a generalized equiangular measurement: $\{P_{\alpha,k}^{A}:\,k=1,\ldots,M_{\alpha}^{A};\,\alpha=1,\ldots,N_{A}\}$ and analogically for $B$. To measure correlations between subsystems, one introduces the correlation matrix with elements
$\mathcal{P}_{\alpha,k;\beta,\ell}=\mathrm{Tr}\Big{[}\rho(P_{\alpha,k}^{A}\otimes P_{\beta,\ell}^{B})\Big{]}.$ (42)
Now, separability conditions follow in terms of its trace or trace norm $\|X\|_{\mathrm{Tr}}=\mathrm{Tr}\sqrt{X^{\dagger}X}$.
Proposition 4.
If a bipartite state $\rho$ is separable, then $\|\mathcal{P}\|_{\mathrm{Tr}}\leq\sqrt{C_{A}C_{B}}$. If additionally $\sum_{\alpha=1}^{N_{A}}M_{\alpha}^{A}=\sum_{\beta=1}^{N_{B}}M_{\beta}^{B}$, then the necessary separability condition reads $\mathrm{Tr}\mathcal{P}\leq\frac{C_{A}+C_{B}}{2}$. Here, $C_{A/B}$ are the indices of coincidence corresponding to conical 2-designs $\{P_{\alpha,k}^{A/B}\}$.
Proof.
For a product state $\rho=\rho_{A}\otimes\rho_{B}$, one has $\mathcal{P}_{\alpha,k;\beta,\ell}=p_{\alpha,k}^{A}p_{\beta,\ell}^{B}$ with $p_{\alpha,k}^{A}=\mathrm{Tr}(P_{\alpha,k}^{A}\rho_{A})$ and $p_{\alpha,k}^{B}=\mathrm{Tr}(P_{\alpha,k}^{B}\rho_{B})$. Therefore, using the methods from ref. ESIC , it follows that
$\|\mathcal{P}\|_{{\rm{tr}}}=\mathrm{Tr}\sqrt{\mathcal{P}^{2}}=\sqrt{\sum_{\alpha=1}^{N_{A}}\sum_{k=1}^{M_{\alpha}^{A}}(p_{\alpha,k}^{A})^{2}}\sqrt{\sum_{\beta=1}^{N_{B}}\sum_{\ell=1}^{M_{\beta}^{B}}(p_{\beta,\ell}^{B})^{2}}=\sqrt{C_{A}C_{B}}.$ (43)
Note that the trace of $\mathcal{P}$ is well defined only under the additional condition that $\mathcal{P}$ is a square matrix. This is the case as long as $\sum_{\alpha=1}^{N_{A}}M_{\alpha}^{A}=\sum_{\beta=1}^{N_{B}}M_{\beta}^{B}$. Then,
$\mathrm{Tr}\mathcal{P}=\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}p_{\alpha,k}^{A}p_{\alpha,k}^{B}\leq\frac{1}{2}\sum_{\alpha=1}^{N}\sum_{k=1}^{M_{\alpha}}\Big{[}(p_{\alpha,k}^{A})^{2}+(p_{\alpha,k}^{B})^{2}\Big{]}=\frac{C_{A}+C_{B}}{2}.$ (44)
Finally, from the convexity of the trace norm and linearity of the trace, the above conditions in terms of $\mathrm{Tr}\mathcal{P}$ and $\|\mathcal{P}\|_{\mathrm{Tr}}$ extend from product states to all separable states. ∎
The above results include as special cases the separability criteria for mutually unbiased bases Spengler and measurements ChenMa , general SIC POVMs ChenLi , $(N,M)$-POVMs SIC-MUB , generalized symmetric measurements SIC-MUB_general , and measurements based on equiangular tight frames W_ETS .
4 Acknowledgements
This research was funded in whole or in part by the National Science Centre, Poland, Grant number 2021/43/D/ST2/00102. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.